OpenStudy (winsteria):

MEDAL AND FAN what is the solution to this proportion? 12/15 = 8/x A. 12 B. 9 C. 10 D. 16

OpenStudy (winsteria):

@acxbox22 @Agl202

OpenStudy (anonymous):

\[\frac{12}{15}=\frac{8}{x}\] \[12x=8\times 15\] \[x=\frac{8\times 15}{12}\]

OpenStudy (winsteria):

8 x 15 = 120

OpenStudy (acxbox22):

yes now divide that by 12 to get x

OpenStudy (anonymous):

in your head of course since \[\frac{120}{12}=10\]

OpenStudy (winsteria):

Thank you

OpenStudy (anonymous):

yw